OpenStudy (itsmichelle29):

What are the period and phase shift for f(x) = −4 tan(x − π)?

OpenStudy (itsmichelle29):

thats all wrong

OpenStudy (anonymous):

The work is correct.

Nnesha (nnesha):

lol

OpenStudy (itsmichelle29):

no because the function she wrote as given was not my function

OpenStudy (anonymous):

Standard equation for tangent function: y=A tan(Bx - C) + D period=π/B, Phase shift=C/B , amplitude |A|, vertical shift D For given function: y= -4 tan (x- π) = -4 tan ( 1*x - π) B=1, C = π period=π/B=π/1 phase shift=C/B=π/1=π (to the right) The Phase Shift is how far the function is horizontally to the right of the parent function y= tan(x)